OpenStudy (anonymous):

Ok, I'm having trouble with a few integral problems, and finding the area. y=x+1, y=9-x^2,x=-1, x=2 I keep getting 109/6 but my book says i should be getting 117/6

OpenStudy (anonymous):

If someone gets another answer plz let me know how

OpenStudy (raden):

yeah, i have got the answer : 117/6=39/2 try to make the picture first, then you will see the border of two curves, next process the integral to find the area

OpenStudy (anonymous):

Ya that is exactly how I did it but I must be doing something wrong.

OpenStudy (anonymous):

for the integral did you get (8x -(x^3)/3-(x^2)/2)?

OpenStudy (raden):

no, for the integral i get (9-x^2 - (x+1) = (-x^2-x+8)

OpenStudy (anonymous):

right you took the integral of that but after the integral was taken what did you get?

OpenStudy (anonymous):

do you see what I'm saying?

OpenStudy (raden):

for the integral (-x^2-x+8), [-1,2] i think u can solve it

OpenStudy (raden):

sorry, i cant say more,, my English too bad :(

OpenStudy (anonymous):

I got -(x^3)/3-(x^2)/2+8x from the integral is that what you got

OpenStudy (raden):

the first, subtitute for x=2, we will get .... next, subtitute for x=-1, we will get .... subtract both

OpenStudy (anonymous):

ya, Ive done that like 20 times still getting 109/6..... even with a calculator

OpenStudy (raden):

u should get the answer : (-8/3 - 2 + 16) - (1/3 - 1/2 -8)

OpenStudy (anonymous):

k, thank you!! I was forgetting the 3 on the bottom of my x^3 term, I knew it had to be something simple i was just missing it

OpenStudy (raden):

yaw, sorry i cant explaind to u too more... i have a problem, with language here :'(